A particular medication dosage is 20.1 mg/kg of body weight. If 1.00 mL of the medication contains 50.0 mg, how many mL of the medication should a child weighing 59.0 lbs. receive?
1
Expert's answer
2019-09-17T04:55:41-0400

1 lbs = 0.453 Kg

Thus 59 lbs = 59 * 0.453 = 26.761 kg

1 kg requires 20.1 mg of does

26.761 kg = 20.1 * 26.761 = 537.89 mg


here 1 mL contains 50 mg

thus 537.89 mg will be contained in = 537.89/50 = 10.75 mL